OpenStudy (anonymous):

What is the measure of one angle in a regular 24-gon?

OpenStudy (chosenmatt):

what do you think?

OpenStudy (anonymous):

I want to say 165

jimthompson5910 (jim_thompson5910):

If you want the interior angle, use the formula I = 180(n-2)/n and you are correct. When n = 24, 180(n-2)/n = 180(24-2)/24 = 165

OpenStudy (anonymous):

Awesome thank you very much for your help! Makes more sense now! ^>^

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Thanks for the metal!